OpenStudy (anonymous):

Convert -630° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed.

OpenStudy (anonymous):

-7pi/2

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Correct! \[(-630*\pi)/180=(-7\pi/2)\]